OpenStudy (anonymous):

if one peice of a yardstick is 8 inches longer than the other how long is each? i need work shown step by step

OpenStudy (anonymous):

a yardstick is 36 inches long if you have a piece that is 8 inches long, the other piece is 36-8 inches long, or 28 inches long

OpenStudy (anonymous):

x+8=36 x+8-8=36-8 x=28

OpenStudy (anonymous):

I think he misread the problem. Let one pice be 8 inches long. Then the other piece is x+8 inches long. And when you put them together... you know how long it should be total. Can you set up the equation now?

OpenStudy (anonymous):

yes i did read it wrong, i thought it said one piece was 8 inches.

OpenStudy (anonymous):

So it would be. x+(x+8)=36 2x+8=36 2x=28 x=14 so one piece is 14 in and the other is 22in

OpenStudy (anonymous):

yep